OpenStudy (firejay5):

Find the difference @SmoothMath 3x/ 3x^2 + 10x -8 - 9x/5x^2 +11x - 36

OpenStudy (firejay5):

@SmoothMath Stay on here for 8 of theses problems

OpenStudy (anonymous):

\(\Large \frac{3x}{3x^2}+10x-8-\frac{9x}{5x^2} +11x-36\) ??

OpenStudy (anonymous):

Or is it \(\Large \frac{3x}{3x^2+10x-8} - \frac{9x}{5x^2+11x-36}\)

OpenStudy (firejay5):

correct, the one with 3x^2 on the bottom

OpenStudy (firejay5):

the second one

OpenStudy (firejay5):

3x3x2+10x−8−9x5x2+11x−36

OpenStudy (anonymous):

Well with the two fractions, you can simplify them. Then you can just group like terms.